### v1

# Large Language Models Might Not Care What You Are Saying: Prompt Format Beats Descriptions

Abstract
With the help of in-context learning (ICL), large language models (LLMs) have achieved impressive performance across various tasks. However, the function of descriptive instructions during ICL remains under-explored. In this work, we propose an ensemble prompt framework to describe the selection criteria of multiple in-context examples, and preliminary experiments on machine translation (MT) across six translation directions confirm that this framework boosts ICL perfromance. But to our surprise, LLMs might not necessarily care what the descriptions actually say, and the performance gain is primarily caused by the ensemble format, since the framework could lead to improvement even with random descriptive nouns. We further apply this new ensemble prompt on a range of commonsense, math, logical reasoning and hallucination tasks with three LLMs and achieve promising results, suggesting again that designing a proper prompt format would be much more effective and efficient than paying effort into specific descriptions. Our code will be publicly available once this paper is published.
1 Introduction
In-context learning (ICL) boosts the performance of large language models (LLMs) across numerous natural language processing (NLP) tasks, where LLMs are presented with in-context examples containing both input and ground truth output (Brown et al. 2020; Dong et al. 2023). Many works have verified the vital role of in-context examples in ICL (Wang et al. 2023; Wei et al. 2023). However, Min et al. (2022) find that ground truth labels might not be the key to ICL performance on classification tasks.
The selection of in-context examples has been proven significant to the performance of ICL (Rubin, Herzig, and Berant 2022) and there have been various works on in-context example selection (Agrawal et al. 2023; Li et al. 2023; Ye et al. 2023). Besides diverse approaches of selecting examples, no existing work has tried to explicitly tell LLMs in what way those specific examples are selected. We hypothesize that if LLMs are prompted with instructions describing the properties of selected in-context examples, they might learn better from these examples, since instruction following is one of LLMs’ most important qualities nowadays (Ouyang et al. 2022; Peng et al. 2023; Zhang et al. 2024).
Recently, Tang, Wang, and Wu (2024) propose to prompt LLMs with examples selected based on both word-level and syntax-level criteria for machine translation (MT) for better ICL performance. This inspires us to tell LLMs where different in-context examples come from when examples are selected by multiple methods.
In our experiments on MT, we first select in-context examples based on lexical and syntactic similarity for each test input separately. Then we combine both to construct the complete set of examples, with half word-level examples and half syntax-level examples. Further, we devise a novel ensemble prompt framework, adding example-level instructions to describe that the following examples are with similar words or similar syntax.
Experimental results on MT demonstrate that adding such ensemble prompt framework does improve LLMs’ performance over conventional prompts. Meanwhile, we find that when the example-level descriptions do not correspond to the source of in-context examples (e.g., examples are said to be with similar syntax when they are actually selected based on word-level similarity) or are completely nonsense (e.g., the descriptive noun ”words” or ”syntax” is replaced with a random noun or the qualifier ”similar” is replaced with ”different”), LLMs still benefit from the prompt. These surprising results indicate that in fact LLMs might not care what the descriptions say and are more sensitive to the prompt format. In other words, a proper format can be much more effective than well-designed descriptions in ICL.
To further verify the superiority of the ensemble framework, we present empirical evaluations on commonsense, math, logical reasoning and hallucination benchmarks (including nine datasets in total) across three small-scale LLMs (Alpaca, Llama2 and Mistral) and one large-scale LLM (GPT-3.5). The novel prompt framework is able to achieve promising results even with the descriptive nouns in the prompt being random nouns, further suggesting that a proper prompt format would be much more effective and efficient compared with laborious design of detailed and specific descriptions.
Our contributions can be summarized as follows:
  • We analyze the effect of prompt descriptions on ICL performance and find that LLMs might not necessarily care what users actually say in descriptions, while they are more sensitive to the prompt format. To the best of our knowledge, this is the first time that the effect of descriptions in prompts is specifically studied.
  • We present a simple yet effective prompt framework that is preliminarily proven feasible on MT through comprehensive experiments across six translation directions. Promising experimental results on three LLMs (Alpaca, Llama2 and Mistral) further verify the superiority of the novel framework on a range of commonsense, math, logical reasoning and hallucination tasks (including nine datasets in total).
2 Prompting LLMs for Machine Translation
Primarily, we focus on MT, a typical generation task that many works have studied ICL on it. Recently, various approaches of selecting in-context examples have been proposed for MT (Agrawal et al. 2023; Kumar et al. 2023; Tang, Wang, and Wu 2024). However, to the best of our knowledge, no previous work has tried to make LLMs aware of in what way those specific in-context examples are selected.
We assume that LLMs would perform better when they are told the reasons for selecting those examples. Tang, Wang, and Wu (2024) select examples based on a combination of word-level and syntax-level criteria, which inspires us to present an ensemble prompt to enable LLMs to clearly know the reasons behind example sampling. In addition, to have a comprehensive idea of whether LLMs really know what is said in the descriptions, we design some prompt variants that are less meaningful or completely nonsense.
2.1 In-context Example Selection for Machine Translation
For word-level examples, we simply select them using BM25 (Bassani 2023). For syntax-level examples, we use the simplified polynomial algorithm proposed by Tang, Wang, and Wu (2024) to convert dependency trees into polynomials and compute syntactic similarity based on the Manhattan distances (Craw 2017) between polynomial terms. For simplicity, we select syntax-level examples based on the Top-$k$ results of the polynomial algorithm instead of the coverage-based approach proposed by Tang, Wang, and Wu (2024). For brevity, we denote the syntax-level algorithm by ”polynomial”.
To combine examples from word-level and syntax-level criteria, we simply concatenate them. For example, the first half of in-context examples are selected by BM25 while the remaining half are selected by the polynomial algorithm.
2.2 A New Ensemble Prompt Framework
To maintain consistency, all our MT experiments use four in-context examples.
First of all, we use the most regular prompt without any example-level descriptions as a baseline (referred to as Vanilla), which is shown in Figure 1.
Figure 1: Template and Alpaca’s example of Vanilla.
In the template, ”Task-level Instruction” instructs the model to do the current task (MT here). ”Example $A_{i}$” and ”Example $B_{i}$” denote the $i$-th example from selection approach $A$ (e.g., BM25) and $B$ (e.g., polynomial) respectively, all containing both source language inputs and target language translations. ”Test Input” refers to the source language input of the test sample, which requires the LLM to translate it into the target language.
Then, we add example-level descriptions for examples from different selection approaches and explicitly instruct the LLM to translate the test input. This prompt framework is referred to as Ensemble and is shown in Figure 2.
Figure 2: Template and Alpaca’s example of Ensemble.
In the template, ”Noun $A$” and ”Noun $B$” describe the examples from selection $A$ and $B$ respectively. For example, the two nouns can be ”words” and ”syntax” to properly describe examples selected by BM25 and polynomial respectively. In this way, we can conveniently control the example-level instructions to tell the LLM why those in-context examples are used.
Figure 3: Main results on XGLM and Alpaca, showing the performance gain of different prompts over the vanilla prompt, averaged over all six translation directions. Each cluster presents the results of a selection of in-context examples and each bar in it presents the result of a prompt. ”Ens.”, ”W.”, ”Syn.”, ”Sem.”, ”Diff.”, ”Rand.”, ”Poly.” refer to ”Ensemble”, ”Word”, ”Syntax”, ”Semantics”, ”Different”, ”Random”, ”Polynomial” respectively.
2.3 Experimental Setup
Language ISO Code Dataset #Pairs (M)
German DE Europarl 1.8
French FR Europarl 1.9
Russian RU ParaCrawl 5.4
Table 1: Data statistics.
Datasets
We perform evaluation on the devtest set of FLORES-101 (Goyal et al. 2022), which contains 1012 sentences with translations in 101 languages. We experiment between English and three common languages: German, French and Russian. We use Europarl (Koehn 2005) for German and French and ParaCrawl (Bañón et al. 2020) for Russian as example database, from which we select in-context examples. Detailed statistics are presented in Table 1.
Evaluation Metrics
We report COMET (Rei et al. 2020) scores from wmt20-comet-da  [^1] https://huggingface.co/Unbabel/wmt20-comet-da, which is considered a superior metric for MT today (Kocmi et al. 2021).
Language Models
[…]
Example Selection
To maintain consistency, all our MT experiments use 4 in-context examples. We evaluate different ways of selecting examples for comparison. Note that if all 4 examples are selected by the same method, the first two are considered examples from $A$ and the last two are considered from $B$ in the Ensemble template in Figure 2.
  • Random: The 4 examples are randomly sampled from the example database. We report the average result of three different random seeds.
  • BM25: We retrieve the top-$4$ matching examples for each test input using BM25 (Bassani 2023).
  • Polynomial: It is rather time-consuming to retrieve examples from databases containing millions of data using the polynomial algorithm. Following Tang, Wang, and Wu (2024), we instead re-rank the top-$100$ examples retrieved by BM25 using polynomial and the top-$4$ are used as final in-context examples.
  • BM25 + Polynomial: To combine examples with both lexical and syntactic similarity, we simply concatenate examples from BM25 and Polynomial. Specifically, the first two examples are from BM25 and the remaining two are from Polynomial.
  • Polynomial + BM25: The first two examples are from Polynomial and the remaining two are from BM25.
Prompts
We design various prompts to explore whether LLMs can benefit from explicit descriptions of examples and whether they really understand the meaning of descriptions.
  • Vanilla: The normal prompt without any example-level descriptions shown in Figure 1.
  • Ensemble (Word + Syntax): The Ensemble prompt shown in Figure 2, with Noun $A$ and Noun $B$ being ”words” and ”syntax” respectively, which semantically corresponds to BM25 + Polynomial examples but is converse to Polynomial + BM25.
  • Ensemble (Syntax + Word): Noun $A$ and Noun $B$ are ”syntax” and ”words” respectively, which semantically matches Polynomial + BM25 examples but mismatches BM25 + Polynomial.
  • Different Ensemble (Word + Syntax): Noun $A$ and Noun $B$ are still ”words” and ”syntax” respectively but the qualifier ”similar” is replaced with ”different”. This prompt aims to find out whether LLMs know the meaning of ”different/similar” and care the semantics of descriptions.
  • Ensemble (Word + Semantics): Noun $A$ and Noun $B$ are ”words” and ”semantics” respectively.
  • Ensemble (Random + Random): For each test input, Noun $A$ and Noun $B$ are different random English nouns sampled using Wonderwords [^2] https://github.com/mrmaxguns/wonderwordsmodule. This aims to further explore LLMs’ understanding of descriptions.
2.4 Main Results
To give a quick view of LLMs’ performance on MT, Table 2 presents the COMET scores of baselines averaged over all six translation directions using Vanilla prompt.
Selection XGLM Alpaca
Random 54.07 55.42
BM25 55.00 56.27
Polynomial 55.52 56.13
BM25 + Polynomial 56.17 56.18
Polynomial + BM25 56.18 55.49
Table 2: Results of Vanilla baselines.
Main results are shown in Figure 3. For convenient comparison, we present the performance gain of different Ensemble prompts over Vanilla with different selections of in-context examples and the results are averaged over all six translation directions.
As can be seen from the results, those ”correct” prompts, exactly corresponding to the selection of in-context examples (e.g., Ensemble (Word + Syntax) with BM25 + Polynomial examples and Ensemble (Syntax + Word) with Polynomial + BM25 examples), do bring some help as expected. However, when the prompt does not correspond to the selection of examples (is not ”correct”), the performance improves as well and sometimes even more than those ”correct” cases. For example, on XGLM with BM25 + Polynomial examples, Ensemble (Syntax + Word) improves more than Ensemble (Word + Syntax), even though the former is completely reversed. On Alpaca with BM25 + Polynomial examples, Ensemble (Word + Semantics) improves more than Ensemble (Word + Syntax), albeit the examples with similar syntax do not necessarily bear similar semantics. More interestingly, Different Ensemble (Word + Syntax), telling the LLM that the in-context examples are with different properties, is able to beat ”correct” prompts sometimes (e.g., on XGLM with BM25 + Polynomial examples and Alpaca with Polynomial + BM25 examples).
Surprisingly, no matter how in-context examples are selected and whether the prompts are ”correct”, Ensemble prompts bring improvement in most cases. Even Ensemble (Random + Random), in which example-level descriptions are with random nouns and could be completely nonsense (like ”examples with similar nobody”), brings improvement in most cases, especially obtaining the most gain on Alpaca with Polynomial + BM25 examples compared with other prompts, correct or incorrect. These results indicate that LLMs might not really take the example-level descriptions into consideration during ICL. In other words, they might not necessarily care what users say in the descriptions.
Compared with proper descriptions, it seems the format of prompts matters more. For example, on Alpaca with Random examples, no matter what the example-level descriptions say, all Ensemble prompts bring nearly equal improvement over Vanilla. This indicates that Ensemble is a superior format compared with Vanilla in this case.
To sum up, the experimental results on MT suggest that a proper prompt format leads to better ICL performance of LLMs while a careful design of descriptions might be less effective.
2.5 Ablation Study
To better understand how the Ensemble format brings improvement, we perform ablation experiments over the organization of the prompt:
  • Ensemble (Random + Random): The Ensemble prompt with random nouns in the example-level descriptions as described in Section 2.3.
  • Single (Random): Organized based on Figure 2, but the second description with Noun $B$ is removed. There is only one example-level description above all the four examples, with Noun $A$ being a random noun.
  • Single (Example): Organized based on Figure 2, but the second description with Noun $B$ is removed. There is only one example-level description above all the four examples, being ”Examples:” only, without any further descriptions. This prompt only informs the LLM that the following four instances are examples and does not describe their properties.
  • Vanilla (Translate): Organized based on Figure 2, but both the two descriptions are removed. The only difference with Vanilla is the translation instruction ”Translate the following sentence:” before the test input. This prompt only informs the LLM to translate the test input and tells nothing about the in-context examples.
Figure 4: Ablation studies over the organization of the prompt, showing the performance gain of different prompts over Vanilla, averaged over all six translation directions. ”Rand.”, ”Poly.”, ”Ens.”, ”Sgl.”, ”V.”, ”Trans.” refer to ”Random”, ”Polynomial”, ”Ensemble”, ”Single”, ”Vanilla”, ”Translate” respectively.
Results are presented in Figure 4, showing that removing one or two example-level descriptions or removing the random noun describing the property of in-context examples hurt the performance gain in most cases. On XGLM, only the original Ensemble format performs better than Vanilla. On Alpaca, Single (Random), Single (Example) and Vanilla (Translate) bring less improvement than Ensemble (Random + Random) in more than half of the cases.
The ablation experiments demonstrate that in MT, our proposed Ensemble is a relatively superior prompt format, obtaining better results than other prompt variants.
2.6 Analysis via Attention Weights
From-To Attention Weight (×1e-4)
Ens. (W. + Syn.) Ens. (Rand. + Rand.)
EA - NA 69.03 60.63
EB - NA 22.49 23.28
EB - NB 16.42 22.93
T - NA 10.41 18.49
T - NB 10.89 18.24
Table 3: Results of attention weights on XGLM, comparing Ensemble (Word + Syntax) and Ensemble (Random + Random). ”EA”, ”EB”, ”NA”, ”NB” and ”T” denote examples from $A$ and $B$, Noun $A$ and Noun $B$, and the target position, respectively.
To have a better idea of the internal mechanism of LLMs when prompted with different prompts, we calculate the attention weights between different prompt parts. We focus on three components: in-context examples (from $A$ or $B$), the target position where the model starts to generate predictions (following Wang et al. (2023), we use the final token in the input) and the two descriptive nouns (Noun $A$ and Noun $B$). We obtain the attention weights averaged over all attention heads from the attention matrix in the last layer. Following Wang et al. (2023) and Zhao et al. (2021), we use the first subword if the noun is tokenized into multiple tokens. All the results are averaged over all six language directions.
Results on XGLM with BM25 + Polynomial examples are presented in Table 3. If the model really cares what the descriptions say, its attention to meaningful descriptive nouns should be much greater than those meaningless. However, the attention weights from examples to descriptive nouns of Ensemble (Word + Syntax) are not necessarily greater than those of Ensemble (Random + Random), indicating that the model does not care what the nouns actually are when managing in-context examples. The weights between the target position and the nouns of the former are even lower than the latter, showing that the model pays little attention to the actual meaning of the descriptions during generation. We conjecture that the relatively higher weights from the target position to random nouns might be due to the confusion of the model caused by those nonsense. To sum up, the attention weights further confirm our claims.
2.7 Discussion
Above results show that LLMs benefit from our Ensemble prompts in most cases. However, the benefit comes from a proper format rather than the carefully designed meaningful descriptions (e.g., ”similar words” and ”similar syntax”). This demonstrates that LLMs might not necessarily care what users say in the descriptions but is more sensitive to the format of prompts. In other words, designing a proper prompt format would be more efficient than paying a lot of effort into looking for a perfect description.
In the following section, we apply the superior Ensemble format to more tasks to further verify its generalization.
3 Generalizing the New Ensemble Prompt Framework to More Tasks
To further verify our conclusion obtained from MT that our proposed Ensemble framework improves ICL even when the example-level descriptions are incorrect or meaningless, we perform the comparison between Vanilla and Ensemble (Random + Random), which we would refer to as ERR, on more types of tasks across different language models.
Figure 5: Results on all nine datasets across three small-scale models. In the ”Date” subplot, the scores of Llama2 and Mistral under the Vanilla prompt are too low to be visible bars in the chart.
3.1 Experimental Setup
Datasets
We conduct a total of nine benchmarks, covering four types of tasks: commonsense QA, logical reasoning, arithmetic reasoning, and hallucination detection.
For commonsense QA, we adopt four datasets that either are widely used or focus on certain specific aspects of knowledge. The widely-used CSQA (Talmor et al. 2019) features commonsense questions about the world involving complex semantics that often require prior knowledge. StrategyQA (Geva et al. 2021) challenges models to infer implicit reasoning steps using a strategy to answer questions. Additionally, we choose two specialized evaluation sets from the BIG-bench (Srivastava et al. 2023): Date Understanding, which asks models to infer the date from a context, and Sports Understanding, which involves assessing the plausibility of sentences related to sports.
For logical reasoning task, we choose Logical Fallacy and Three Objects (a subset of Logical Deduction) from Big-bench (Srivastava et al. 2023). Logical Fallacy aims to test the model’s ability to identify whether there are fallacies in a given logical reasoning, and Three Objects requires the model to infer the order of a sequence of objects from a set of minimal conditions.
To explore the performance of ERR on math word problems, we adopt the following two datasets: GSM8K (Cobbe et al. 2021), which consists of high quality free-response grade school math problems, and AQuA (Ling et al. 2017), containing the algebraic word problems in the form of multiple-choice questions.
In addition, in order to explore whether ERR could alleviate LLMs’ hallucination, we choose Known Unknowns from Big-bench (Srivastava et al. 2023).
The details of splitting training set (example database) and test set are presented in supplementary materials.
Evaluation Metric
These nine datasets are either in the form of multiple-choice questions or free-response questions with standard answers, so we use accuracy as the metric for all of them.
Language Models
We evaluate three open source LLMs with 7B parameters, including Alpaca (Taori et al. 2023), Llama2 (Touvron et al. 2023b), and Mistral (Jiang et al. 2023). To assess the effect of ERR on more powerful models, we also evaluate GPT-3.5 (Ouyang et al. 2022). We use Llama-2-7b-chat-hf, Mistral-7B-Instruct-v0.2 and gpt-3.5-turbo-0125333https://openai.com/api/ for Llama2, Mistral and GPT-3.5 respectively.
Example Selection
Note that randomly selected examples combined with ERR have already brought non-trivial improvements to MT. Therefore, for each dataset discussed in this section, we randomly select a uniform set of examples (4-shot) for all test inputs without applying any carefully designed selection method, in order to focus on and verify the simple yet effective and universal nature of ERR.
Prompts
We compare ERR with Vanilla across different datasets and LLMs. Given that these tasks usually involve reasoning, on which chain-of-thought (CoT) is commonly utilized (Wei et al. 2022), we experiment both without CoT (”w/o CoT”, which are identical to the original templates) and with CoT (”w/ CoT”). This allows us to examine both the orthogonality and compatibility with CoT of ERR , as well as assess its performance across various models and tasks. Specifically, we evaluate Vanilla (w/o CoT), Vanilla (w/ CoT), ERR (w/o CoT), and ERR (w/ CoT). Due to space constraints, examples of prompt templates discussed in this section are provided in the supplementary materials.
Figure 6: Results of all the four types of tasks on GPT-3.5.
3.2 Results of Small-scale Models
Results across all nine datasets and three small-scale models (Alpaca, Llama2 and Mistral) are illustrated in Figure 5.
The results demonstrate that ERR (w/ CoT), achieved by integrating CoT with our proposed prompt framework, either significantly outperforms or matches Vanilla (w/ CoT) in 25 out of 27 experiments (covering 9 datasets and 3 models). The exceptions are Alpaca on the Sports dataset and Mistral on the AQuA dataset, where ERR (w/ CoT) shows somewhat lower performance compared to Vanilla (w/ CoT). When CoT is not incorporated, ERR generally performs much better than or on par with Vanilla, except for the Logical Fallacy and Sports datasets with Llama2. In these cases, Llama2 frequently produces emojis or alternative questions under the ERR framework.
Surprisingly, ERR (w/o CoT) sometimes even surpasses Vanilla (w/ CoT), suggesting that the ERR framework alone can offer more improvements than CoT. This highlights the value of ERR and reaffirms that the format plays a crucial role in enhancing the model’s problem-solving capabilities.
In terms of models, the performance of ERR on Alpaca is far less impressive than on Llama2 and Mistral. This phenomenon also appears in GPT-3.5, which will be discussed in the next section. Another noteworthy observation is that while ERR outperforms Vanilla with Llama2 across most datasets, its performance on the Logical Fallacy and Sports datasets is notably poor. As mentioned earlier, Llama2 almost always responds with confused emojis for Logical Fallacy questions and outputs questions like ”plausible or implausible?” for Sports, leading to predominantly incorrect answers. Further investigation into these issues is left for future work.
In summary, without using any carefully designed selection methods, directly filling the randomly selected examples into the ERR framework brings significant improvement to various reasoning tasks and even alleviates the hallucination of models in most cases, no matter how meaningless and incorrect the example-level descriptions are. Moreover, ERR can work perfectly with CoT. Therefore, at least for relatively small models, this simple but effective trick is worth introducing into prompt engineering for various tasks.
3.3 Results of GPT-3.5
As shown in Figure 6, ERR performs similarly to Vanilla across every datasets using GPT-3.5. Although the ERR format does not bring significant improvement to these tasks with GPT-3.5 and Alpaca (as shown in Figure 5), the fact remains that the incorrect or meaningless example-level descriptions caused by random nouns do not have much negative impact on GPT-3.5, a sufficiently powerful model, or Alpaca, which has strong instruction-following capabilities. In some cases, it even slightly improves performance (e.g., ERR (w/ CoT) outperforms Vanilla (w/ CoT) on AQuA and Known Unknowns datasets). In other words, LLMs might not necessarily care what users actually say to describe the provided examples while they are more sensitive to the format of prompts, which is in line with our findings obtained from MT.
3.4 Discussion
Based on the experiments conducted on both small-scale and large-scale models, we can conclude that ERR is a simple yet practical and universal prompt framework. It can enhance problem-solving capabilities in small models and be applied to large models without the risk of performance degradation due to the meaningless noise within it. In other words, there might be less need to meticulously select examples or design detailed descriptions. Instead, you can uniformly and efficiently apply ERR to prompt engineering of various tasks with different models.
4 Related Work
In-context Example Selection
Rubin, Herzig, and Berant (2022) suggest that LLMs’ ICL performance strongly depends on the selection of in-context examples. In consequence, many works have been trying to explore ways of selecting better in-context examples in recent years. Li et al. (2023) train a unified in-context example retriever across a wide range of tasks. Ye et al. (2023) select examples based on both relevance and diversity, with the help of determinantal point processes. Agrawal et al. (2023) ensure n-gram coverage to select better examples for MT. Kumar et al. (2023) train an in-context example scorer for MT based on several features. Tang, Wang, and Wu (2024) propose to combine both word-level and syntax-level coverage when selecting examples for MT.
Mechanism of In-context Learning
With the popularity of ICL, there have been numerous studies on analyzing the mechanism of ICL. One stream of these studies focuses on explaining the essence of ICL, relating ICL to gradient descent (Von Oswald et al. 2023), implicit Bayesian inference (Xie et al. 2022), induction heads completing token sequences based on similar context (Olsson et al. 2022), generation maintaining coherency (Sia and Duh 2023), creation of task vectors based on in-context examples (Hendel, Geva, and Globerson 2023), etc. The other stream focuses on the role of in-context examples, especially labels of these examples. Min et al. (2022) find that ground truth labels are not necessary and LLMs perform fairly well even with random labels, suggesting that LLMs mainly learn the label space, the input distribution and the overall task format from in-context examples. Wang et al. (2023) find that label words play the role of anchors that aggregating information of the whole examples and serve as a reference for LLMs’ final predictions. Wei et al. (2023) find that larger language models are able to override semantic priors and successfully learn from in-context examples with flipped labels or even semantically-unrelated labels.
5 Conclusion
In this work, we analyze the effect of descriptive instructions in prompts during ICL and propose an ensemble prompt framework describing the properties of in-context examples selected by different selection methods. Experimental results on MT indicate that while LLMs are sensitive to prompt formats, they might not necessarily care the actual meaning of the descriptions and the framework improves LLMs’ performance even with meaningless descriptions compared with the conventional prompt. We further apply the ensemble framework to four other NLP tasks and find that it achieves promising results, especially on small-scale models. These results suggest that rather than working hard on well-designed descriptions, making use of a proper prompt format would be more effective and efficient.