### v4

# Large Language Models Might Not Care What You Are Saying: Prompt Format Beats Descriptions

Abstract
With the help of in-context learning (ICL), large language models (LLMs) have achieved impressive performance across various tasks. However, the function of descriptive instructions during ICL remains under-explored. In this work, we propose an ensemble prompt framework to describe the selection criteria of multiple in-context examples, and preliminary experiments on machine translation (MT) across six translation directions confirm that this framework boosts ICL performance. But to our surprise, LLMs might not care what the descriptions actually say, and the performance gain is primarily caused by the ensemble format, since it could lead to improvement even with random descriptive nouns. We further apply this new ensemble framework on a range of commonsense, math, logical reasoning and hallucination tasks with three LLMs and achieve promising results, suggesting again that designing a proper prompt format would be much more effective and efficient than paying effort into specific descriptions. Our code will be publicly available once this paper is published.
Large Language Models Might Not Care What You Are Saying:
Prompt Format Beats Descriptions
Chenming Tang  Zhixiang Wang  Hao Sun  Yunfang Wuthanks:    Corresponding author. National Key Laboratory for Multimedia Information Processing, Peking University MOE Key Laboratory of Computational Linguistics, Peking University School of Computer Science, Peking University {tangchenming, ekko}@stu.pku.edu.cn{2301213218, wuyf}@pku.edu.cn
1 Introduction
In-context learning (ICL) boosts the performance of large language models (LLMs) across numerous natural language processing (NLP) tasks, where LLMs are presented with in-context examples containing input and ground truth output Brown et al. (2020); Dong et al. (2023). Many works have verified the vital role of in-context examples in ICL Wang et al. (2023); Wei et al. (2023). However, Min et al. Min et al. (2022) find that ground truth labels might not be the key to ICL performance on classification tasks.
The selection of in-context examples has been proven significant to the performance of ICL Rubin et al. (2022) and there have been various works on in-context example selection Agrawal et al. (2023); Li et al. (2023); Ye et al. (2023). Besides diverse approaches of selecting examples, no existing work has tried to explicitly tell LLMs in what way those specific examples are selected. We hypothesize that if LLMs are prompted with instructions describing the properties of selected in-context examples, they might learn better from these examples, since instruction following is one of LLMs’ most important qualities nowadays Ouyang et al. (2022); Peng et al. (2023); Zhang et al. (2024).
Recently, Tang et al. (2024) prompt LLMs with examples selected based on both word-level and syntax-level criteria for machine translation (MT) for better ICL performance. This inspires us to tell LLMs where different in-context examples come from when they are selected by multiple methods.
Figure 1: Template and Alpaca’s example of Ensemble.
In our experiments on MT, we first select in-context examples based on lexical and syntactic similarity for each test input separately. Then we combine both to construct the complete set of examples, with half word-level examples and half syntax-level examples. Further, we devise a novel ensemble prompt framework (as shown in the left part "Prompt Template" of Figure 1), adding example-level instructions to describe that the following examples are with similar words or similar syntax.
Experimental results on MT demonstrate that adding such ensemble prompt framework does improve LLMs’ performance over conventional prompts. Meanwhile, we find that when the example-level descriptions do not correspond to the source of in-context examples or are completely nonsense, LLMs still benefit from the prompt. These surprising results indicate that in fact LLMs might not care what the descriptions say and are more sensitive to the prompt format. In other words, a proper format can be much more effective than well-designed descriptions in ICL.
To further verify the superiority of the ensemble framework, we present empirical evaluations on commonsense, math, logical reasoning and hallucination benchmarks (including nine datasets in total) across three small-scale LLMs (Alpaca, Llama3 and Mistral) and one large-scale LLM (GPT-3.5). The novel prompt framework is able to achieve promising results even with the descriptive nouns in the prompt being random nouns, further suggesting that a proper prompt format would be much more effective and efficient compared with laborious design of detailed and specific descriptions.
There are a few studies very related to our work. Min et al. (2022) find that the labels of in-context examples do not need to be correct for classification tasks. Wei et al. (2023) find that larger language models successfully learn from in-context examples even when the labels are flipped or semantically-unrelated. Our work is different from the above in that we focus on the meaning of descriptions rather than labels in ICL and our finding is that the format of prompts is more important than carefully designed descriptions.
Our contributions can be summarized as follows:
  • For the first time, we specifically analyze the effect of prompt descriptions on ICL performance and find that LLMs might not care what users actually say in descriptions, while they are more sensitive to the prompt format.
  • We present a simple yet effective prompt framework that is proven feasible on MT through comprehensive experiments across six translation directions. Promising experimental results on three LLMs further verify the superiority of the novel framework on a range of commonsense, math, logical reasoning and hallucination tasks.
2 Prompting LLMs for MT
Primarily, we focus on MT, a typical generation task. Recently, various approaches of selecting in-context examples have been proposed for MT Agrawal et al. (2023); Kumar et al. (2023); Tang et al. (2024). However, no existing work has tried to make LLMs aware of in what way those specific in-context examples are selected.
We assume that LLMs would perform better when they are told the reasons for selecting those examples. Tang et al. Tang et al. (2024) select examples based on a combination of word-level and syntax-level criteria, which inspires us to present an ensemble prompt framework to make LLMs clearly know the reasons behind example selection. In addition, to have a comprehensive idea of whether LLMs really know what is said in the descriptions, we design some prompt variants that are less meaningful or completely nonsense.
2.1 In-context Example Selection for MT
For word-level examples, we simply select them using BM25 Bassani (2023). For syntax-level examples, we use the top-$k$ polynomial algorithm proposed by Tang et al. Tang et al. (2024) to convert dependency trees into polynomials and compute syntactic similarity based on the Manhattan distances Craw (2017) between polynomial terms. For brevity, we denote the syntax-level algorithm by "Polynomial".
To combine word-level and syntax-level examples, we simply concatenate them. For example, the first and the remaining half of examples are selected by BM25 and Polynomial respectively.
2.2 A New Ensemble Prompt Framework
To maintain consistency, all our MT experiments use four in-context examples.
First of all, we use the most regular prompt without any example-level descriptions as baseline (referred to as Vanilla), which is shown in Figure 2.
Figure 2: Template and Alpaca’s example of Vanilla.
In the template, "Task-level Instruction" instructs the model to do the current task (MT here). "Example $A_{i}$" and "Example $B_{i}$" denote the $i$-th example from selection approach $A$ (e.g., BM25) and $B$ (e.g., Polynomial) respectively, all containing both source language inputs and target language translations. "Test Input" refers to the source language input of the test sample, which requires the LLM to translate it into the target language.
Then, we add example-level descriptions for examples from different selection approaches and explicitly instruct the LLM to translate the test input. This prompt framework is referred to as Ensemble and is shown in Figure 1 as presented in Section 1. "Noun $A$" and "Noun $B$" describe the examples from selection $A$ and $B$ respectively. For example, the two nouns can be "words" and "syntax" to properly describe examples selected by BM25 and Polynomial respectively. In this way, we can conveniently control the example-level descriptions to tell the LLM why those examples are used.
Figure 3: Main results on XGLM and Alpaca, showing the performance gain of different prompts over the vanilla prompt, averaged over all six translation directions. Each cluster presents the results of a selection of in-context examples and each bar in it presents the result of a prompt. "Ens.", "W.", "Syn.", "Sem.", "Diff.", "Rand.", "Poly." refer to "Ensemble", "Word", "Syntax", "Semantics", "Different", "Random", "Polynomial" respectively.
2.3 Experimental Setup
Language ISO Code Dataset #Pairs (M)
German DE Europarl 1.8
French FR Europarl 1.9
Russian RU ParaCrawl 5.4
Table 1: Data statistics.
2.3.1 Datasets
We perform evaluation on the devtest set of FLORES-101 Goyal et al. (2022), which contains 1012 sentences with translations in 101 languages. We experiment between English and three common languages: German, French and Russian. We use Europarl Koehn (2005) for German and French and ParaCrawl Bañón et al. (2020) for Russian as example database, from which we select in-context examples. Detailed statistics are in Table 1.
2.3.2 Evaluation Metrics
We report COMET Rei et al. (2020) scores from wmt20-comet-da  [^1] https://huggingface.co/Unbabel/wmt20-comet-da, which is considered a superior metric for MT today Kocmi et al. (2021).
2.3.3 Language Models
[…]
2.3.4 Example Selection
To maintain consistency, all our MT experiments use 4 in-context examples. We evaluate different ways of selecting examples for comparison. Note that if all 4 examples are selected by the same method, the first two are considered examples from $A$ and the last two are considered from $B$ in the Ensemble template in Figure 1.
Random: The 4 examples are randomly sampled from the example database. We report the average result of three different random seeds.
BM25: We retrieve the top-$4$ matching examples for each test input using BM25 Bassani (2023).
Polynomial: It is rather time-consuming to retrieve examples from databases containing millions of data using the Polynomial algorithm. Following Tang et al. Tang et al. (2024), we instead re-rank the top-$100$ examples retrieved by BM25 using Polynomial and the top-$4$ are used as final in-context examples.
BM25 + Polynomial: To combine examples with both lexical and syntactic similarity, we simply concatenate examples from BM25 and Polynomial. Specifically, the first two examples are from BM25 and the remaining two are from Polynomial.
Polynomial + BM25: The first two examples are from Polynomial and the remaining two are from BM25.
2.3.5 Prompts
We design various prompts to explore whether LLMs can benefit from explicit descriptions of examples and whether they really understand the meaning of descriptions.
Vanilla: The normal prompt without any example-level descriptions shown in Figure 2.
Ensemble (Word + Syntax): The Ensemble prompt shown in Figure 1, with Noun $A$ and Noun $B$ being "words" and "syntax" respectively, which semantically corresponds to BM25 + Polynomial examples but is converse to Polynomial + BM25.
Ensemble (Syntax + Word): Noun $A$ and Noun $B$ are "syntax" and "words" respectively, which semantically matches Polynomial + BM25 examples but mismatches BM25 + Polynomial.
Different Ensemble (Word + Syntax): Noun $A$ and Noun $B$ are still "words" and "syntax" respectively but the qualifier "similar" is replaced with "different". This prompt aims to find out whether LLMs know the meaning of "different/similar" and care the semantics of descriptions.
Ensemble (Word + Semantics): Noun $A$ and Noun $B$ are "words" and "semantics" respectively.
Ensemble (Random + Random): For each input, Noun $A$ and Noun $B$ are different random English nouns sampled using Wonderwords [^2] https://github.com/mrmaxguns/wonderwordsmodule, aiming to explore LLMs’ understanding of descriptions.
Detailed templates and examples of the above prompts are shown in Appendix A.1.
2.4 Main Results
To give a quick view of LLMs’ MT performance, Table 2 shows the COMET scores of Vanilla baselines averaged over six translation directions.
Selection XGLM Alpaca
Random 54.07 55.42
BM25 55.00 56.27
Polynomial 55.52 56.13
BM25 + Polynomial 56.17 56.18
Polynomial + BM25 56.18 55.49
Table 2: Results of Vanilla baselines.
Main results are shown in Figure 3. For convenient comparison, we present the performance gain of different Ensemble prompts over Vanilla with different selections of in-context examples and the results are averaged over six translation directions.
As can be seen from the results, those "correct" prompts, exactly corresponding to the selection of in-context examples (e.g., Ensemble (Word + Syntax) with BM25 + Polynomial examples and Ensemble (Syntax + Word) with Polynomial + BM25 examples), do bring some help as expected. However, when the prompt does not correspond to the selection of examples (is "incorrect"), the performance improves as well and sometimes even more than those "correct" cases. For example, on XGLM with BM25 + Polynomial examples, Ensemble (Syntax + Word) improves more than Ensemble (Word + Syntax), even though the former is completely reversed. On Alpaca with BM25 + Polynomial examples, Ensemble (Word + Semantics) improves more than Ensemble (Word + Syntax), albeit the examples with similar syntax do not necessarily bear similar semantics. More interestingly, Different Ensemble (Word + Syntax), telling the LLM that the in-context examples are with different properties, is able to beat "correct" prompts sometimes (e.g., on XGLM with BM25 + Polynomial examples and Alpaca with Polynomial + BM25 examples).
Surprisingly, no matter how in-context examples are selected and whether the prompts are "correct", Ensemble prompts bring improvement in most cases. Even Ensemble (Random + Random), in which example-level descriptions are with random nouns and could be completely nonsense (like "examples with similar nobody"), brings improvement in most cases, especially obtaining the most gain on Alpaca with Polynomial + BM25 examples compared with other prompts, correct or incorrect. These results indicate that LLMs might not really take the example-level descriptions into consideration during ICL. In other words, they might not necessarily care what users say in the descriptions.
Compared with proper descriptions, it seems the format of prompts matters more. For example, on Alpaca with Random examples, no matter what the example-level descriptions say, all Ensemble prompts bring nearly equal improvement over Vanilla. This indicates that Ensemble is a superior format compared with Vanilla in this case.
To sum up, the experimental results on MT suggest that a proper prompt format leads to better ICL performance of LLMs while a careful design of descriptions might be less effective.
2.5 Ablation Study
To better understand how the Ensemble format brings improvement, we perform ablation experiments over the organization of the prompt:
Ensemble (Random + Random): The Ensemble prompt with random nouns in the example-level descriptions as described in Section 2.3.
Single (Random): Organized based on Figure 1, but the second description is removed. There is only one example-level description above the four examples, where Noun $A$ is a random noun.
Single (Example): Organized based on Figure 1, but the second description is removed. There is only one example-level description above the four examples, being "Examples:" only, without any further descriptions. This prompt only informs the LLM that the following four instances are examples and does not describe their properties.
Vanilla (Translate): Organized based on Figure 1, but both the two descriptions are removed. The only difference with Vanilla is the translation instruction "Translate the following sentence:" before the test input. This prompt only informs the LLM to translate the test input and tells nothing about the in-context examples.
Detailed templates and examples of the above prompts are shown in Appendix A.2.
Figure 4: Ablation studies over the organization of the prompt, showing the performance gain of different prompts over Vanilla, averaged over all six translation directions. "Rand.", "Poly.", "Ens.", "Sgl.", "V.", "Trans." refer to "Random", "Polynomial", "Ensemble", "Single", "Vanilla", "Translate" respectively.
Figure 5: Attention weights (×1e-4) on XGLM of all 32 layers with BM25 + Polynomial examples. EWS and ERR denotes Ensemble (Word + Syntax) and Ensemble (Random + Random) respectively.
Results are presented in Figure 4, showing that removing one or two example-level descriptions or removing the random noun describing the property of in-context examples hurt the performance gain in most cases. On XGLM, only the original Ensemble format performs better than Vanilla. On Alpaca, Single (Random), Single (Example) and Vanilla (Translate) bring less improvement than Ensemble (Random + Random) in more than half of the cases.
Ablation experiments suggest that in MT, our proposed Ensemble is a relatively superior prompt format, performing better than other variants.
2.6 Analysis via Attention Weights
To have a better idea of the internal mechanism of LLMs when prompted with different prompts, we calculate the attention weights between different prompt components. We focus on three components: in-context examples (from $A$ or $B$, denoted by "Example-A" and "Example-B"), the target position (denoted by "Target") where the model starts to generate predictions (following Wang et al. Wang et al. (2023), we use the final token in the input) and the two descriptive nouns ("Noun-A" and "Noun-B"). We obtain the attention weights averaged over all attention heads from the attention matrix across all the layers. All the results are averaged over all six language directions.
Results comparing Ensemble (Word + Syntax) (EWS) and Ensemble (Random + Random) (ERR) on XGLM with BM25 + Polynomial examples are presented in Figure 5 (for results on Alpaca, refer to Appendix B). If the model really cares what the descriptions say, its attention to meaningful descriptive nouns (in EWS) should be much greater than those meaningless (in ERR). However, in most cases, EWS performs no higher than ERR, indicating that the model does not really care what the descriptive nouns actually are. "Target to Noun-A" is a special case, where EWS is high in shallow layers. But in deeper layers, EWS falls behind and ERR takes the lead. This shows that the model might pay more attention to the meaningful noun when understanding the context in shallow layers but gradually forgets it when it comes to generation in deeper layers. In a word, the attention weights further confirm our claim that LLMs do not really care what the descriptive nouns are in most cases.
2.7 Discussion
Above results show that LLMs benefit from our Ensemble prompts in most cases. However, the benefit comes from a proper format rather than the meaningful descriptions (e.g., "similar words" and "similar syntax"). This demonstrates that LLMs might not care what users say in the descriptions but is more sensitive to the format of prompts. In other words, designing a proper prompt format would be more efficient than paying a lot of effort into looking for a perfect description.
In the next section, we apply Ensemble format to more tasks to further verify its generalizability.
3 Generalizing the New Ensemble Prompt Framework to More Tasks
To further verify our conclusion obtained from MT that our proposed Ensemble framework improves ICL even when the example-level descriptions are incorrect or meaningless, we perform the comparison between Vanilla and Ensemble (Random + Random), which we would refer to as ERR, on more types of tasks across different language models.
Figure 6: Results on nine datasets across three small-scale models. In the "Date" subplot, the score of Mistral under the Vanilla prompt is too low to be a visible bar in the chart.
3.1 Experimental Setup
3.1.1 Datasets
We use a total of nine benchmarks, covering four task types: commonsense QA, logical reasoning, arithmetic reasoning, and hallucination detection.
For commonsense QA, we adopt four datasets. The widely-used CSQA Talmor et al. (2019) features commonsense questions about the world involving complex semantics requiring prior knowledge. StrategyQA Geva et al. (2021) challenges models to infer implicit reasoning steps using a strategy to answer questions. We also choose two specialized evaluation sets from BIG-bench Srivastava et al. (2023): Date Understanding, which asks models to infer the date from a context, and Sports Understanding, which involves assessing the plausibility of sentences related to sports.
For logical reasoning task, we choose Logical Fallacy and Three Objects (a subset of Logical Deduction) from Big-bench Srivastava et al. (2023). Logical Fallacy aims to test the model’s ability to identify whether there are fallacies in a given logical reasoning, and Three Objects requires the model to infer the order of a sequence of objects from a set of minimal conditions.
To explore the performance of ERR on math word problems, we adopt the following two datasets: GSM8K Cobbe et al. (2021), which consists of high quality free-response grade school math problems, and AQuA Ling et al. (2017), containing the algebraic word problems in the form of multiple-choice questions.
In addition, to explore whether ERR could alleviate LLMs’ hallucination, we choose Known Unknowns from Big-bench Srivastava et al. (2023).
Dataset Test Inputs
CSQA 1221
StrategyQA 1012
Date 365
Sports 996
Logical Fallacy 1012
Three Objects 296
Known Unknowns 42
GSM8K 1319
AQuA 254
Table 3: Number of test inputs for each dataset.
The number of test inputs for each dataset is listed in Table 3. Details of splitting training set (example database) and test set are in Appendix C.
3.1.2 Evaluation Metric
These nine datasets are either in the form of multiple-choice questions or free-response questions with standard answers, so we use accuracy as the metric for all of them.
3.1.3 Language Models
We evaluate three open source LLMs with around 7B parameters, including Alpaca Taori et al. (2023), Llama3 Grattafiori et al. (2024), and Mistral Jiang et al. (2023). To assess the effect of ERR on more powerful models, we also evaluate GPT-3.5 Ouyang et al. (2022). We use Llama-3.1-8B, Mistral-7B-Instruct-v0.2 and gpt-3.5-turbo-0125333https://openai.com/api/ for Llama3, Mistral and GPT-3.5 respectively.
3.1.4 Example Selection
Note that randomly selected examples combined with ERR have already brought non-trivial improvements to MT. Therefore, for each dataset discussed in this section, we randomly select a uniform set of examples (4-shot) for all test inputs without applying any carefully designed selection method, in order to focus on and verify the simple yet effective and universal nature of ERR.
3.1.5 Prompts
We compare ERR with Vanilla across different datasets and LLMs. Given that these tasks usually involve reasoning, on which chain-of-thought (CoT) is commonly utilized Wei et al. (2022), we experiment both without CoT ("w/o CoT", which are identical to the original templates) and with CoT ("w/ CoT"). This allows us to examine both the orthogonality and compatibility with CoT of ERR , as well as assess its performance across various models and tasks. Specifically, we evaluate Vanilla (w/o CoT), Vanilla (w/ CoT), ERR (w/o CoT), and ERR (w/ CoT). Due to space constraints, examples of prompt templates discussed in this section are provided in Appendix D.
Figure 7: Results of the four types of tasks on GPT-3.5.
3.2 Results of Small-scale Models
Results across all nine datasets and three small-scale models (Alpaca, Llama3 and Mistral) are illustrated in Figure 6.
The results demonstrate that ERR (w/ CoT), achieved by integrating CoT with our proposed prompt framework, either significantly outperforms or matches Vanilla (w/ CoT) in 25 out of 27 experiments (covering 9 datasets and 3 models). The exceptions are Alpaca on the Sports dataset and Mistral on the AQuA dataset, where ERR (w/ CoT) shows somewhat lower performance compared to Vanilla (w/ CoT). When CoT is not incorporated, ERR generally performs much better than or on par with Vanilla, except for the Sports dataset with Llama3, where ERR performs a little poorer.
Surprisingly, ERR (w/o CoT) sometimes even surpasses Vanilla (w/ CoT), suggesting that the ERR framework alone can offer more improvements than CoT. This highlights the value of ERR and reaffirms that the format plays a crucial role in enhancing LLMs’ problem-solving capabilities. In terms of models, the performance of ERR on Alpaca is far less impressive than on Llama3 and Mistral, which may be because Alpaca has strong instruction-following capabilities and is more robust to different prompts.
In summary, without using any carefully designed selection methods, directly filling the randomly selected examples into the ERR framework brings significant improvement to various reasoning tasks and even alleviates the hallucination of models in most cases, no matter how meaningless and incorrect the example-level descriptions are. Moreover, ERR can work perfectly with CoT. Therefore, at least for relatively small models, this simple but effective trick is worth introducing into prompt engineering for various tasks.
We also experiment with Llama2 Touvron et al. (2023b) and the results are in Appendix E. The overall trend is consistent with Llama3.
3.3 Results of GPT-3.5
As shown in Figure 7, ERR performs similarly to Vanilla across every dataset using GPT-3.5. Although the ERR format does not bring significant improvement to these tasks with GPT-3.5 and Alpaca (as shown in Figure 6), the fact remains that the incorrect or meaningless example-level descriptions caused by random nouns do not have much negative impact on GPT-3.5, a sufficiently powerful model, or Alpaca, which has strong instruction-following capabilities. In some cases, it even slightly improves performance (e.g., ERR (w/ CoT) outperforms Vanilla (w/ CoT) on AQuA and Known Unknowns). In other words, LLMs might not care what users actually say to describe the provided examples while they are more sensitive to the format of prompts, which is in line with our findings obtained from MT.
3.4 Discussion
Based on the experiments conducted on both small-scale and large-scale models, we can conclude that ERR is a simple yet practical and universal prompt framework. It can enhance problem-solving capabilities in small models and be applied to large models without the risk of performance degradation due to the meaningless noise within it. In other words, there might be less need to meticulously select examples or design detailed descriptions. Instead, you can uniformly and efficiently apply ERR to various tasks with different models.
4 Related Work
In-context Example Selection
Rubin et al. Rubin et al. (2022) suggest that LLMs’ ICL performance strongly depends on the selection of in-context examples. In consequence, many works have been trying to explore ways of selecting better in-context examples in recent years. Li et al. Li et al. (2023) train a unified in-context example retriever across a wide range of tasks. Ye et al. Ye et al. (2023) select examples based on both relevance and diversity, with the help of determinantal point processes. Agrawal et al. Agrawal et al. (2023) ensure n-gram coverage to select better examples for MT. Kumar et al. Kumar et al. (2023) train an in-context example scorer for MT based on several features. Tang et al. Tang et al. (2024) propose to combine both word-level and syntax-level coverage when selecting examples for MT.
Mechanism of In-context Learning
With the popularity of ICL, there have been numerous studies on analyzing the mechanism of ICL. One stream of these studies focuses on explaining the essence of ICL, relating ICL to gradient descent Von Oswald et al. (2023), implicit Bayesian inference Xie et al. (2022), induction heads completing token sequences based on similar context Olsson et al. (2022), generation maintaining coherency Sia and Duh (2023), creation of task vectors based on in-context examples Hendel et al. (2023), etc. The other stream focuses on the role of in-context examples, especially labels of these examples. Min et al. Min et al. (2022) find that ground truth labels are not necessary and LLMs perform fairly well even with random labels. Wang et al. Wang et al. (2023) find that label words play the role of anchors that aggregating information of the whole examples and serve as a reference for LLMs’ final predictions. Wei et al. Wei et al. (2023) find that larger language models can override semantic priors and learn from in-context examples with flipped labels or semantically-unrelated labels.
5 Conclusion
In this work, we analyze the effect of descriptive instructions in prompts during ICL and propose an Ensemble prompt framework describing the properties of in-context examples selected by different methods. Experimental results on MT indicate that while LLMs are sensitive to prompt formats, they might not care the actual meaning of the descriptions and the framework improves LLMs’ performance even with meaningless descriptions compared with the conventional prompt. We further apply the Ensemble framework to four other NLP tasks and find that it achieves promising results, especially on small-scale models. These results suggest that rather than working hard on well-designed descriptions, making use of a proper prompt format would be more effective and efficient.
Limitations
First, since there are so many open-source LLMs in the world nowadays, it is impossible to experiment with all existing models and thus our work only employ several commonly-used LLMs. Second, since we do not have access to the pre-training or post-training process of LLMs (either open-source or close-source), our analysis of the mechanism of ICL could be somewhat superficial. The behavior of LLMs can be highly subject to their training data, which we have no access to. Lastly, although we reveal that ERR is a superior prompt format for several models, it could still be a local optimum and how to effectively search for a best prompt format for different models and tasks is still under-explored, which we leave for future work.
Appendix A Prompt Templates of MT Experiments
Figure 8: Template and example of Vanilla.
Figure 9: Template and example of Ensemble (Word + Syntax).
Figure 10: Template and example of Ensemble (Syntax + Word).
Figure 11: Template and example of Different Ensemble (Word + Syntax).
Figure 12: Template and example of Ensemble (Word + Semantics).
Figure 13: Template and example of Ensemble (Random + Random).
Figure 14: Template and example of Single (Random).
Figure 15: Template and example of Single (Example).
Figure 16: Template and example of Vanilla (Translate).
A.1 Main Experiments
Templates and prompt examples of Vanilla, Ensemble (Word + Syntax), Ensemble (Syntax + Word), Different Ensemble (Word + Syntax), Ensemble (Word + Semantics) and Ensemble (Random + Random) are shown in Figure 8-13.
A.2 Ablations
Templates and prompt examples of Ensemble (Random + Random), Single (Random), Single (Example), Vanilla (Translate) are shown in Figure 13-16.
Appendix B Attention Weights on Alpaca
Figure 17: Attention weights (×1e-4) on Alpaca of all 32 layers with BM25 + Polynomial examples. EWS and ERR denotes Ensemble (Word + Syntax) and Ensemble (Random + Random) respectively.
Figure 17 presents the attention weights on Alpaca. For example-to-noun attention weights, ERR is close to EWS. For target-to-noun attention weights, EWS is higher in shallow layers but falls behind ERR in deeper layers, especially in the last layer. This demonstrates that Alpaca might pay more attention to the meaningful words ("word" and "syntax") when understanding the context in shallow layers but gradually forgets them when it comes to generation in the deeper layers. In short, EWS performs no higher than ERR in most cases.
Appendix C Dataset Details for Reasoning Tasks
We list the details of splitting training set (example database) and test set for our conducted reasoning tasks, covering four types and nine datasets. We set random seed for all possible shuffling and sampling operations to $42$. Note that we experiment with 4-shot for all datasets.
C.1 Datasets Fetched from Exclusive Source
  • CSQA Talmor et al. (2019): https://www.tau-nlp.org/commonsenseqa. We follow the official split and select the training set as our example database and the dev set as our test set. Because the training set itself is randomly divided from the whole dataset, we directly select examples from it in order.
  • GSM8K Cobbe et al. (2021): https://github.com/openai/grade-school-math. We select the test.jsonl as our test set and the train.jsonl as our example database and randomly sample four examples from it.
  • AQuA Ling et al. (2017): https://github.com/google-deepmind/AQuA. We select the test.json as our test set. Since the original training set is relatively large, for simplicity, we directly copy the four examples listed in the supplementary materials of Wei et al. (2022) and we ensure that these four examples do not appear in the test set.
C.2 Datasets Fetched from The Big-bench
For StrategyQA Geva et al. (2021), Date, Sports, Logical Fallacy, Three Objects, and Known Unknowns, we fetched them from the Big-bench Srivastava et al. (2023). Each of them has a task.json. We randomly shuffle the task.json and split it to a training set and a test set. Then we select examples from the training set in order.
Specifically, the principle for splitting the training and test sets is as follows: If the total number of samples exceeds 1,012 a lot, we retain 1,012 samples as the test set and use the remainder as the training set. Otherwise, we select four examples for the training set and use the rest for testing. For the Sports and Logical Fallacy datasets, which have only two possible answers (similar to binary classification), we first separate the positive and negative examples, shuffle them individually, and then construct the test set and training set. The test set is composed of an equal number of positive and negative examples, with the remaining samples used as the training set.
Appendix D Prompts for Reasoning Tasks Used in this Work
Figure 18: Prompt for CSQA.
Figure 19: Prompt for StrategyQA.
Figure 20: Prompt for Date.
Figure 21: Prompt for Sports.
Figure 22: Prompt for Logical Fallacy.
Figure 23: Prompt for Three Objects.
Figure 24: Prompt for Known Unknowns.
Figure 25: Prompt for GSM8K. For this dataset, we let LLMs first generate reasoning and then answer under the "w/ CoT" setting.
Figure 26: Prompt for AQuA.
Figure 18-26 show the examples of ERR (w/ CoT) prompt for respective datasets. Some tasks contain Answer Choices. In order to save space, the blank lines between the options are replaced with spaces in those figures. Each Figure has grey text for reasoning, cyan text for the parts of ERR that are unique to Vanilla, and italic words in the cyan text representing random nouns. Therefore, deleting the grey text gives ERR (w/o CoT), keeping the grey text but deleting the cyan text gives Vanilla (w/ CoT), and deleting both the cyan and grey text gives Vanilla (w/o CoT). [^4] When changing ”w/ CoT” to ”w/o CoT”, you may also need to replace ”So the answer is …” with ”The answer is …” for syntactical reasons. The reasoning is generated by ChatGPT  [^5] https://chatgpt.com/. Note the ChatGPT is not the same as GPT-3.5 we used for experiments.
Appendix E Results of Llama2
Figure 27: Results of Llama2 on the nine datasets.
Results of Llama2-7B-chat-hf Touvron et al. (2023b) on the nine datasets are presented in Figure 27. While ERR outperforms Vanilla with Llama2 across most datasets, its performance on Logical Fallacy and Sports is notably poor. Llama2 almost always responds with confused emojis for Logical Fallacy questions and outputs questions like "plausible or implausible?" for Sports, leading to predominantly incorrect answers. Further investigation into these issues is left for future work.
Appendix F Computational Details
F.1 Hardware
Inference of LLMs runs on an NVIDIA A40 GPU (with memory of 48 GB). Other experiments run on Intel® Xeon® Gold 6348 CPU (with memory of 256 GB).
F.2 Software
Our OS: Ubuntu 20.04.6 LTS. Our code: Python only. Libraries and packages are specified in the source code.
Appendix G Licenses
Artifact License
XGLM MIT
Alpaca Apache-2.0
Llama Llama Community License Agreement
Mistral Apache-2.0
COMET Apache-2.0
FLORES-101 CC-BY-SA-4.0
Europarl Unknown
ParaCrawl CC0
CSQA CC-BY-SA-4.0
StrategyQA MIT
BIG-bench Apache-2.0
GSM8K MIT
AQuA Apache-2.0
Table 4: Licenses of scientific artifacts we use.
The licenses of the scientific artifacts we use are shown in Table 4.